# Capturing Polysemanticity with PRISM: A Multi-Concept Feature Description Framework

### v3

@@ A.1 Related Work
A wide variety of works seek to examine the mechanisms of generative AI systems. Many initially focused on attributions in input space [8, 9], with later approaches focused on discovering circuit and graph structures [3, 4, 5], as well as self-explanations generated by the model [10, 11]. To extract relevant representations and model features, a variety of works have explored interpretability for analyzing concepts  [73], representations [74], and clustering [75] in the context of attributions of sentence summaries [76] and concept discovery in LLMs [77]. In the following, we will focus specifically on studies and methods centered on feature descriptions.
@@ A.2 Feature Description Methods
One of the earliest works on feature description in language models is SASC (Summarize and Score) [78], which generates natural language descriptions of neurons in a pre-trained BERT model. Shortly thereafter, an automated interpretability method for describing all neurons in GPT-2 XL was proposed [21]. This approach analyzes the textual patterns that cause a neuron to activate, and uses GPT-4 as explainer model to generate a description of the neuron’s function. Given a set of token-activation pairs derived from text excerpts and corresponding neuron activations, the explainer model identifies common patterns, based on which it generates a textual description of the neuron’s role. This method has since been widely adopted and further developed, forming the basis for many subsequent methods targeting both individual neurons [23, 25] and SAE features [26, 19, 22, 28, 27, 44, 29, 25]. An overview of representative feature description methods is provided in Table 2.
@@
SASC [78] BERT neuron
GPT-Explain [21] GPT-2 XL neuron
Pythia SAE [26] Pythia-70M and Pythia 410-M SAE feature
Anthropic SAE [19] one-layer transformer SAE feature
GPT-2 SAE [22] GPT-2 Small SAE feature
GPT-4 SAE [22] GPT-4 SAE feature
EleutherAI SAE [28] Llama 3.1 7B & Gemma 2 9B SAE feature
Transluce-Explain [23] Llama 3.1-8B Instruct neuron
Llama Scope [27] Llama 3.1 8B Base SAE feature
Gemma Scope [44] Gemma 2 SAE feature
Goodfire SAE [29] Llama 3.3 70B SAE feature
Output-Centric neuron [25] Llama 3.1 8B Instruct neuron
Output-Centric SAE [25] Gemma 2.2B, GPT-2 Small, Llama 3.1 8B SAE feature
@@ A.3 Description Scoring Details
Figure 7 illustrates the CoSy evaluation procedure [39] as adapted for language models. As the control dataset $\mathbb{X}_{0}$, we use a subset of 1,000 randomly sampled entries from Cosmopedia [40]. For each candidate description of a target feature, we use Gemini 1.5 Pro [46] to generate 10 concept-specific text samples, each with a maximum length of 512 tokens. These samples form the concept dataset $\mathbb{X}_{1}$. The generation prompt is shown in Figure 8. We then pass both datasets through the model to extract activations corresponding to the target feature. We then use Average Pooling as aggregation function $\sigma:\mathbb{R}^{d}\rightarrow\mathbb{R}$ to each activation vector to obtain scalar representations:
@@ A.4 Benchmark Experiment Details
We use the publicly available GPT-Explain descriptions [^7] https://github.com/openai/automated-interpretability/tree/main?tab=readme-ov-file#public-datasets and the Output-Centric feature descriptions [^8] https://github.com/yoavgur/Feature-Descriptions as comparison.
@@
AUC and MAD Distributions
To better understand the standard deviation observed in our benchmark results, we provide distribution plots of the evaluation metrics. Figure 13(a) shows the distribution of AUC scores across all evaluated model features, while Figure 13(b) presents the distribution of MAD scores.
(a) Distributions of PRISM (max) AUC scores across different models and layers.
@@ A.5 Extended Ablation Analysis
Table 3: Impact of the number of clusters ($k$) on feature description quality for GPT-2 XL. A cluster size of $k=5$ corresponds to the original implementation used throughout the paper. Increasing $k$ improves best-case descriptions (higher PRISM max scores) by capturing more specific activation patterns, but lowers average interpretability (lower PRISM mean scores) as patterns fragment across clusters. This illustrates the tradeoff between precision and redundancy. AUC values are reported with 95% confidence intervals; MAD values are reported with standard deviations.
@@
AUC ($\uparrow$) MAD ($\uparrow$) AUC ($\uparrow$) MAD ($\uparrow$)
[…]
@@
To test the robustness of our framework across different text generators, we extended the description generation experiments beyond Gemini 1.5 Pro to several open-source language models: Qwen3 32B [48] [^13] https://huggingface.co/Qwen/Qwen3-32B, Phi-4 [49] [^14] https://huggingface.co/microsoft/phi-4, DeepSeek R1 [50] [^15] https://huggingface.co/deepseek-ai/DeepSeek-R1. These models were used to generate feature descriptions for GPT-2 XL using the same procedure as in our original setup described in Section 4.1. As shown in Table 4, Qwen3 32B achieves performance comparable to Gemini 1.5 Pro, demonstrating that our framework does not depend solely on a single model. While Phi-4 and DeepSeek R1 show slightly lower scores, they still follow the same qualitative trends, demonstrating that our method is robust and generalizes effectively across a range of language models, including accessible open-source options.
Table 4: Ablation study on text generators used for description generation. We report PRISM scores (max and mean) for different LLMs using the original experimental setup. Qwen3 32B achieves performance comparable to Gemini 1.5 Pro, while Phi-4 and DeepSeek R1 exhibit slightly lower scores but maintain the same qualitative trends, demonstrating the framework’s effectiveness across multiple language models. Reported AUC includes 95% confidence intervals; reported MAD includes standard deviations.
@@
AUC ($\uparrow$) MAD ($\uparrow$) AUC ($\uparrow$) MAD ($\uparrow$)
[…]
@@
We evaluated the robustness of our feature descriptions by using Qwen3 32B [48], Phi-4 [49], DeepSeek R1 [50] as alternative text generators in the evaluation step (Appendix A.3 provides details on description scoring). In this process, a set of 10 concept-specific text samples for each feature is generated with the LLM, and AUC and MAD scores are computed to assess description quality for both PRISM and GPT-Explain (Table 5). Although absolute scores are slightly lower when Gemini 1.5 Pro is not used for evaluation, the relative ranking of methods remains consistent. PRISM consistently outperforms GPT-Explain, demonstrating the generalizability of our framework and the robustness of our evaluation setup.
Table 5: Ablation study on text generators used in the evaluation step. We report AUC and MAD scores for PRISM and GPT-Explain when using Gemini 1.5 Pro (original setup) or alternative LLMs (Qwen3 32B, Phi-4, DeepSeek R1) for description scoring. While absolute scores decrease with alternative models, relative rankings remain stable, with PRISM consistently outperforming GPT-Explain. AUC is reported with 95% confidence intervals, and MAD is reported with standard deviations.
@@
AUC ($\uparrow$) MAD ($\uparrow$) AUC ($\uparrow$) MAD ($\uparrow$) AUC ($\uparrow$) MAD ($\uparrow$)
[…]
@@ A.6 Extended Sanity Check Analysis
To ensure that the generated feature descriptions are faithful to the tokens and corresponding text samples in the same cluster, we perform a fully randomized counter probe. For this sanity check, we replace the percentile sampling in Step 1 of Figure 2 with random sampling, drawing random sentences and their corresponding activations from the validation set of the C4 corpus [41]. Since choosing highly activating samples across our random set would bias the random uniform sample distribution, we apply the clustering procedure across all random samples. While we maintain the cluster size, all text samples and highlights are shuffled and assigned to random clusters instead. Both the assignment of random clusters and highlights limits sample similarity. Lastly, we proceed with labeling the randomized clusters. Each cluster consists of unrelated texts and highlights, thus, the LLM should be unable to generate a coherent, shared feature description. As shown in Table6, AUC and MAD scores decrease, confirming that meaningful descriptions depend on the proper grouping of highly activating samples.
@@
We probe whether our feature descriptions truly capture the concepts of their assigned clusters by performing an additional randomization of descriptions. We reassign feature descriptions by randomly sampling from existing descriptions generated for other features within the same model and collect the associated activations. This tests whether explanation quality decreases when descriptions no longer correspond to their clusters, without generating new descriptions. To this end, we use the same feature and layer selection as described in Section 4.1 for GPT-2 XL, which serves as the baseline for non-randomized PRISM performance (see Table 1 GPT-2 XL, MLP neuron). We expect AUC and MAD scores to drop significantly, since the descriptions do not align with the cluster. Table 6 shows that AUC and MAD scores drop significantly compared to the baseline, demonstrating that aligned descriptions are essential for high-quality feature description.
Table 6: Sanity check results for GPT-2 XL. We compare AUC and MAD scores under two randomized settings: (1) Random Sentences, where clusters are formed from unrelated text samples, and (2) Random Descriptions, where existing descriptions are randomly reassigned across features. Both conditions show a notable drop in AUC and MAD scores compared to the Baseline, supporting the faithfulness of our descriptions. AUC reports $95\%$ confidence intervals; MAD shows standard deviations.
@@
AUC ($\uparrow$) MAD ($\uparrow$) AUC ($\uparrow$) MAD ($\uparrow$)
[…]
@@
To verify that polysemanticity scores reflect meaningful semantic relationships rather than artifacts of the embedding process, we perform a sanity check using randomly assigned descriptions. For each feature, five random descriptions (distinct from the true set) are embedded using the same embedding model (gte-Qwen2-1.5B-instruct sentence transformer [45]), and pairwise cosine similarities are computed between them. This is repeated across all features within each layer, and the resulting scores are visualized as box plots grouped by layer (see Figure 17(a)). The difference between true and random scores is generally quite distinct, with random descriptions yielding notably lower similarity values, indicating less semantic coherence. An exception is Llama 3.1 8B Instruct, where cosine similarity scores remain consistently low across layers and frequently approach the random baseline. This indicates that the descriptions associated with Llama’s features are likely more diverse and less semantically aligned compared to other models.
@@
We examine whether feature description quality varies across the percentile activation distribution by segmenting the distribution into quartile-based intervals. For consistency, we apply the same PRISM parameter settings as in the main benchmarking experiments, varying only the start and end points of the percentile range. The baseline setting uses the top 1% of activations (0.99-1.0). Table 7 reports AUC and MAD scores across intervals. As expected, features in the top 25% (0.75-1.0) closely match baseline performance, while lower intervals exhibit reduced scores, indicating a positive correlation between activation strength and description quality.
Table 7: Performance across percentile intervals of feature scores on GPT-2 XL. AUC measures classification performance (95% CI in parentheses), and MAD quantifies activation differences (mean $\pm$ std). Results are shown for both PRISM (max) and PRISM (mean).
@@
AUC ($\uparrow$) MAD ($\uparrow$) AUC ($\uparrow$) MAD ($\uparrow$)
[…]
@@
To further illustrate that sampling from diverse activation ranges is meaningful in practice, we present a selection of feature descriptions. Table 8 shows cases where clusters with diverse mean activation ranges (MeanAct) still achieve high description scores (AUC) (also see Figure 1).
@@
Table 8: A selection of feature descriptions obtained from clusters with diverse mean activation ranges (MeanAct). Despite variation in activation ranges, the descriptions achieve consistently high scores (AUC), illustrating that percentile sampling captures multiple meaningful patterns per feature.
Model, layer, feature Description MeanAct AUC

### v4

@@ A.1 Related Work
A wide variety of works seek to examine the mechanisms of generative AI systems. Many initially focused on attributions in input space [abnar-zuidema-2020-quantifying, ali2022xai], with later approaches focused on discovering circuit and graph structures [wang2023interpretability, nanda2023progress, olahchris-2020-ZoomIntroductionCircuits], as well as self-explanations generated by the model [NEURIPS2018_4c7a167b, madsen-2024-are-self-explanations-faithful]. To extract relevant representations and model features, a variety of works have explored interpretability for analyzing concepts  [chormai24], representations [representations2022], and clustering [kauffmann2022clustering] in the context of attributions of sentence summaries [vasileiou-2024-explaining] and concept discovery in LLMs [poche_consim_2025]. In the following, we will focus specifically on studies and methods centered on feature descriptions.
@@ A.2 Feature Description Methods
One of the earliest works on feature description in language models is SASC (Summarize and Score) [singh-2023-ExplainingBlackBox], which generates natural language descriptions of neurons in a pre-trained BERT model. Shortly thereafter, an automated interpretability method for describing all neurons in GPT-2 XL was proposed [bills-2023-LanguageModelsCan]. This approach analyzes the textual patterns that cause a neuron to activate, and uses GPT-4 as explainer model to generate a description of the neuron’s function. Given a set of token-activation pairs derived from text excerpts and corresponding neuron activations, the explainer model identifies common patterns, based on which it generates a textual description of the neuron’s role. This method has since been widely adopted and further developed, forming the basis for many subsequent methods targeting both individual neurons [choi_scaling_2024, gur-arieh_enhancing_2025] and SAE features [cunningham-2023-SparseAutoencodersFind, bricken-2023-MonosemanticityDecomposingLanguage, neuronpedia, gao_scaling_2024, paulo_automatically_2025, he-2024-LlamaScopeExtracting, lieberum-2024-GemmaScopeOpen, mcgrath--MappingLatentSpace, gur-arieh_enhancing_2025]. An overview of representative feature description methods is provided in Table 2.
@@
SASC [singh-2023-ExplainingBlackBox] BERT neuron
GPT-Explain [bills-2023-LanguageModelsCan] GPT-2 XL neuron
Pythia SAE [cunningham-2023-SparseAutoencodersFind] Pythia 70M and Pythia 410M SAE feature
Anthropic SAE [bricken-2023-MonosemanticityDecomposingLanguage] one-layer transformer SAE feature
Neuronpedia [neuronpedia] DeepSeek R1 Dist Llama 8B, Gemma 2 2B, Gemma 2 2B IT, SAE feature
Gemma 2 9B, Gemma 2 9B IT, GPT OSS 20B,
GPT-2 Small, Llama 3.1 8B, Llama 3.1 8B Instruct,
Pythia 70M Deduped, Qwen 2.5 7B IT, Qwen 3 4B
GPT-2 SAE [gao_scaling_2024] GPT-2 Small SAE feature
GPT-4 SAE [gao_scaling_2024] GPT-4 SAE feature
EleutherAI SAE [paulo_automatically_2025] Llama 3.1 7B & Gemma 2 9B SAE feature
Transluce-Explain [choi_scaling_2024] Llama 3.1 8B Instruct neuron
Llama Scope [he-2024-LlamaScopeExtracting] Llama 3.1 8B Base SAE feature
Gemma Scope [lieberum-2024-GemmaScopeOpen] Gemma 2 SAE feature
Goodfire SAE [mcgrath--MappingLatentSpace] Llama 3.3 70B SAE feature
Output-Centric neuron [gur-arieh_enhancing_2025] Llama 3.1 8B Instruct neuron
Output-Centric SAE [gur-arieh_enhancing_2025] Gemma 2 2B, GPT-2 Small, Llama 3.1 8B SAE feature
@@ A.3 Description Scoring Details
Figure 7 illustrates the CoSy evaluation procedure [kopf-2024-CoSyEvaluatingTextual] as adapted for language models. As the control dataset $\mathbb{X}_{0}$, we use a subset of 1,000 randomly sampled entries from Cosmopedia [benallalloubna-2024-Cosmopedia]. For each candidate description of a target feature, we use Gemini 1.5 Pro [gemini-team-2024-gemini-1-5] to generate 10 concept-specific text samples, each with a maximum length of 512 tokens. These samples form the concept dataset $\mathbb{X}_{1}$. The generation prompt is shown in Figure 8. We then pass both datasets through the model to extract activations corresponding to the target feature. We then use Average Pooling as aggregation function $\sigma:\mathbb{R}^{d}\rightarrow\mathbb{R}$ to each activation vector to obtain scalar representations:
@@ A.4 Benchmark Experiment Details
AUROC and MAD Distributions
To better understand the standard deviation observed in our benchmark results, we provide distribution plots of the evaluation metrics. Figure 13(a) shows the distribution of AUROC scores across all evaluated model features, while Figure 13(b) presents the distribution of MAD scores.
(a) Distributions of PRISM (max) AUROC scores across different models and layers.
@@
Output-Centric Evaluation
In addition to our activation-based metrics, AUC and MAD, we evaluate an output-centric metric, Faithfulness, which quantifies the causal influence of a discovered concept on the model’s output [puri-etal-2025-fade, gur-arieh_enhancing_2025]. Faithfulness measures the causal effect of a feature on model outputs, specifically testing whether directly manipulating a feature’s activations can steer the model to generate content that more strongly reflects the corresponding concept. Following the FADE [^17] https://github.com/brunibrun/FADE implementation of [puri-etal-2025-fade], we compute Faithfulness scores of GPT-2 XL feature descriptions using GPT-4o mini [hurst2024gpt] as the language model and a subset of 10,000 samples from the training split of the Cosmopedia dataset [benallalloubna-2024-Cosmopedia]. To obtain scores for all features, we remove the threshold used for selecting output-centric samples. Figure 13 presents the results of the output-centric evaluation. Across layer groups, PRISM (max) consistently outperforms GPT-Explain, exhibiting trends that closely mirror those observed in the activation-centric evaluation (see Figure LABEL:fig:evaluation-scores), further validating its effectiveness across multiple interpretability criteria.
Figure 13: Output-Centric evaluation of GPT-2 XL across different feature description methods (GPT-Explain, PRISM (max), PRISM (mean)). Each method is evaluated using the output-centric Faithfulness metric, which quantifies the causal influence of a discovered concept on the model’s output. PRISM (max) consistently achieves higher Faithfulness compared to GPT-Explain.
@@ A.5 Extended Ablation Analysis
Table 3: Impact of the number of clusters ($k$) on feature description quality for GPT-2 XL. A cluster size of $k=5$ corresponds to the original implementation used throughout the paper. Increasing $k$ improves best-case descriptions (higher PRISM max scores) by capturing more specific activation patterns, but lowers average interpretability (lower PRISM mean scores) as patterns fragment across clusters. This illustrates the tradeoff between precision and redundancy. AUROC values are reported with 95% confidence intervals; MAD values are reported as the percentage of positive MAD scores.
@@
AUROC ($\uparrow$) MAD ($\uparrow$) AUROC ($\uparrow$) MAD ($\uparrow$)
5 (original) 0.85 (0.78-0.91) 91.67% 0.65 (0.61-0.69) 66.33%
1 0.75 (0.66-0.83) 80.00% 0.75 (0.66-0.83) 80.00%
3 0.82 (0.74-0.89) 81.67% 0.69 (0.64-0.74) 68.33%
10 0.88 (0.83-0.93) 93.33% 0.61 (0.58-0.64) 59.67%
@@
To test the robustness of our framework across different text generators, we extended the description generation experiments beyond Gemini 1.5 Pro to several open-source language models: Qwen3 32B [yang2025qwen3] [^18] https://huggingface.co/Qwen/Qwen3-32B, Phi-4 [abdin2024phi] [^19] https://huggingface.co/microsoft/phi-4, DeepSeek R1 [guo2025deepseek] [^20] https://huggingface.co/deepseek-ai/DeepSeek-R1. These models were used to generate feature descriptions for GPT-2 XL using the same procedure as in our original setup described in Section 4.1. As shown in Table 4, Qwen3 32B achieves performance comparable to Gemini 1.5 Pro, demonstrating that our framework does not depend solely on a single model. While Phi-4 and DeepSeek R1 show slightly lower scores, they still follow the same qualitative trends, demonstrating that our method is robust and generalizes effectively across a range of language models, including accessible open-source options.
Table 4: Ablation study on text generators used for description generation. We report PRISM scores (max and mean) for different LLMs using the original experimental setup. Qwen3 32B achieves performance comparable to Gemini 1.5 Pro, while Phi-4 and DeepSeek R1 exhibit slightly lower scores but maintain the same qualitative trends, demonstrating the framework’s effectiveness across multiple language models. Reported AUROC includes 95% confidence intervals; reported MAD represents the percentage of positive MAD scores.
@@
AUROC ($\uparrow$) MAD ($\uparrow$) AUROC ($\uparrow$) MAD ($\uparrow$)
Gemini 1.5 Pro [gemini-team-2024-gemini-1-5] (original) 0.85 (0.78-0.91) 91.67% 0.65 (0.61-0.69) 66.33%
Qwen3 32B [yang2025qwen3] 0.85 (0.78-0.91) 90.00% 0.65 (0.61-0.69) 64.33%
Phi-4 [abdin2024phi] 0.82 (0.75-0.89) 85.00% 0.61 (0.57-0.65) 61.33%
DeepSeek R1 [guo2025deepseek] 0.79 (0.71-0.87) 78.33% 0.61 (0.56-0.65) 60.33%
@@
We evaluated the robustness of our feature descriptions by using Qwen3 32B [yang2025qwen3], Phi-4 [abdin2024phi], DeepSeek R1 [guo2025deepseek] as alternative text generators in the evaluation step (Appendix A.3 provides details on description scoring). In this process, a set of 10 concept-specific text samples for each feature is generated with the LLM, and AUROC and MAD scores are computed to assess description quality for both PRISM and GPT-Explain (Table 5). Although absolute scores are slightly lower when Gemini 1.5 Pro is not used for evaluation, the relative ranking of methods remains consistent. PRISM consistently outperforms GPT-Explain, demonstrating the generalizability of our framework and the robustness of our evaluation setup.
Table 5: Ablation study on text generators used in the evaluation step. We report AUROC and MAD scores for PRISM and GPT-Explain when using Gemini 1.5 Pro (original setup) or alternative LLMs (Qwen3 32B, Phi-4, DeepSeek R1) for description scoring. While absolute scores decrease with alternative models, relative rankings remain stable, with PRISM consistently outperforming GPT-Explain. AUROC is reported with 95% confidence intervals, and MAD is reported as the percentage of positive MAD scores.
@@
AUROC ($\uparrow$) MAD ($\uparrow$) AUROC ($\uparrow$) MAD ($\uparrow$) AUROC ($\uparrow$) MAD ($\uparrow$)
Gemini 1.5 Pro [gemini-team-2024-gemini-1-5] (original) 0.85 (0.78-0.91) 91.67% 0.65 (0.61-0.69) 66.33% 0.64 (0.56-0.73) 65.00%
Qwen3 32B [yang2025qwen3] 0.58 (0.46-0.69) 58.33% 0.53 (0.48-0.58) 54.00% 0.54 (0.42-0.65) 53.33%
Phi-4 [abdin2024phi] 0.61 (0.50-0.72) 58.33% 0.54 (0.49-0.59) 53.67% 0.56 (0.44-0.67) 58.33%
DeepSeek R1 [guo2025deepseek] 0.71 (0.61-0.80) 73.33% 0.57 (0.52-0.62) 57.67% 0.60 (0.50-0.70) 63.33%
@@ A.6 Extended Sanity Check Analysis
To ensure that the generated feature descriptions are faithful to the tokens and corresponding text samples in the same cluster, we perform a fully randomized counter probe. For this sanity check, we replace the percentile sampling in Step 1 of Figure 2 with random sampling, drawing random sentences and their corresponding activations from the validation set of the C4 corpus [raffel-2020-ExploringLimitsTransfer]. Since choosing highly activating samples across our random set would bias the random uniform sample distribution, we apply the clustering procedure across all random samples. While we maintain the cluster size, all text samples and highlights are shuffled and assigned to random clusters instead. Both the assignment of random clusters and highlights limits sample similarity. Lastly, we proceed with labeling the randomized clusters. Each cluster consists of unrelated texts and highlights, thus, the LLM should be unable to generate a coherent, shared feature description. As shown in Table6, AUROC and MAD scores decrease, confirming that meaningful descriptions depend on the proper grouping of highly activating samples.
@@
We probe whether our feature descriptions truly capture the concepts of their assigned clusters by performing an additional randomization of descriptions. We reassign feature descriptions by randomly sampling from existing descriptions generated for other features within the same model and collect the associated activations. This tests whether explanation quality decreases when descriptions no longer correspond to their clusters, without generating new descriptions. To this end, we use the same feature and layer selection as described in Section 4.1 for GPT-2 XL, which serves as the baseline for non-randomized PRISM performance (see Table 1 GPT-2 XL, MLP neuron). We expect AUROC and MAD scores to drop significantly, since the descriptions do not align with the cluster. Table 6 shows that AUROC and MAD scores drop significantly compared to the baseline, demonstrating that aligned descriptions are essential for high-quality feature description.
Table 6: Sanity check results for GPT-2 XL. We compare AUROC and MAD scores under two randomized settings: (1) Random Sentences, where clusters are formed from unrelated text samples, and (2) Random Descriptions, where existing descriptions are randomly reassigned across features. Both conditions show a notable drop in AUROC and MAD scores compared to the Baseline, supporting the faithfulness of our descriptions. AUROC reports $95\%$ confidence intervals; MAD shows the percentage of positive MAD scores.
@@
AUROC ($\uparrow$) MAD ($\uparrow$) AUROC ($\uparrow$) MAD ($\uparrow$)
Baseline 0.85 (0.78-0.91) 91.67% 0.65 (0.61-0.69) 66.33%
(1) Random Sentences 0.68 (0.59-0.76) 65.00% 0.54 (0.50-0.58) 49.67%
(2) Random Descriptions 0.65 (0.56-0.74) 66.67% 0.52 (0.48-0.57) 49.33%
@@
To verify that polysemanticity scores reflect meaningful semantic relationships rather than artifacts of the embedding process, we perform a sanity check using randomly assigned descriptions. For each feature, five random descriptions (distinct from the true set) are embedded using the same embedding model (gte-Qwen2-1.5B-instruct sentence transformer [li-2023-GeneralTextEmbeddings]), and pairwise cosine similarities are computed between them. This is repeated across all features within each layer, and the resulting scores are visualized as box plots grouped by layer (see Figure 17(a)). The difference between true and random scores is generally quite distinct, with random descriptions yielding notably lower similarity values, indicating less semantic coherence. An exception is Llama 3.1 8B Instruct, where cosine similarity scores remain consistently low across layers and frequently approach the random baseline. This indicates that the descriptions associated with Llama’s features are likely more diverse and less semantically aligned compared to other models.
@@
We examine whether feature description quality varies across the percentile activation distribution by segmenting the distribution into quartile-based intervals. For consistency, we apply the same PRISM parameter settings as in the main benchmarking experiments, varying only the start and end points of the percentile range. The baseline setting uses the top 1% of activations (0.99-1.0). Table 7 reports AUROC and MAD scores across intervals. As expected, features in the top 25% (0.75-1.0) closely match baseline performance, while lower intervals exhibit reduced scores, indicating a positive correlation between activation strength and description quality.
Table 7: Performance across percentile intervals of feature scores on GPT-2 XL. AUROC measures classification performance (95% CI in parentheses), and MAD quantifies activation differences (percentage of positive MAD scores). Results are shown for both PRISM (max) and PRISM (mean).
@@
AUROC ($\uparrow$) MAD ($\uparrow$) AUROC ($\uparrow$) MAD ($\uparrow$)
Baseline 0.85 (0.78-0.91) 91.67% 0.65 (0.61-0.69) 66.33%
0.0 to 0.25 0.69 (0.61-0.77) 66.67% 0.53 (0.49-0.57) 49.33%
0.25 to 0.5 0.74 (0.66-0.82) 70.00% 0.56 (0.52-0.61) 52.67%
0.5 to 0.75 0.71 (0.63-0.79) 73.33% 0.55 (0.51-0.60) 54.33%
0.75 to 1.0 0.85 (0.79-0.91) 90.00% 0.65 (0.61-0.69) 66.33%
@@
To further illustrate that sampling from diverse activation ranges is meaningful in practice, we present a selection of feature descriptions. Table 8 shows cases where clusters with diverse mean activation ranges (MeanAct) still achieve high description scores (AUROC) (also see Figure 1).
@@
Table 8: A selection of feature descriptions obtained from clusters with diverse mean activation ranges (MeanAct). Despite variation in activation ranges, the descriptions achieve consistently high scores (AUROC), illustrating that percentile sampling captures multiple meaningful patterns per feature.
Model, layer, feature Description MeanAct AUROC